3. Band Ltd issued debentures with a face value of N$1 000, which mature in exactly five years from today’s date. The latest interest payment has just been received. The interest is payable at 10% per annum on the face value of the debentures, on a quarterly basis. The market yield for the debentures is quoted as 12% per annum.

The value of a Band Ltd debenture is:

a) N$939
b) N$993
c) N$1 076
d) N$1 000
1
Expert's answer
2020-05-03T17:03:43-0400

We solve by the formula:


"P=\frac{100}{0.12}({1-\frac{1}{(1+0.03)^{20}}})+\frac{1000}{(1+0.12)^5}"


"P=\frac{100}{0.12}({1-\frac{1}{(1+0.03)^{20}}})+\frac{1000}{(1+0.12)^5}"


P=371.94+567.43=939.37


a)answer